Find a power series representation for the function and determine the radius of convergence.

f(x)=((x)/(2-x))^3
Answer #1

we have(1+x) = 1+nx +121x2+ +1n consider = x + (-3 (-3-1)(_). (-3 (-3-1)(-3-2)(-x), for-< 1 レ(1-(-3)(2)+(-3)(-4)(2)-(-3)(4)